2. A steel clip angles are commonly used to transfer loads from horizontal girders to vertical columns. If the reaction of the girders upon the angle is a downward force of 45kN and if two 22mm diameter rivets resist this force, find the average shearing in each of the rivets. Assume that the rivet fills the hole.


1
Expert's answer
2021-09-28T00:57:43-0400

For 2 rivets, there will be 2 upward shearing forces acting as a reaction to the applied force=45kN

By equilibrium of forces;

"\sum F_y=0"

V+V-45=0

V=22.5kN=22.5*103N

Shear stress is given by;

"\tau =\frac{V}{A}"

"\tau =\frac{V}{\frac{π}{4}d^{2}}"

"\tau =\frac{22.5*10^{3}}{\frac{π}{4}*22^{2}}"

=59MPa